General Meeting Summary 6/28/17
Jump to navigation
Jump to search
DETECTORS AT JLAB
- Aerogel Cherenkov Detectors (Arthur, Vardan, ...)
- Setting up two data acquisition systems, one for NPS prototype tests in TEDF clean room, the other in EEL for HMS aerogel Cherenkov detector testing
- communicate list of available components and what is still needed to Steve/Brad - TH send email, Arthur follow up
- Working on SP-11 tray cosmic data taking
- S2 hodoscope plane not fully functional, not in trigger - temporarily setup single scintillator for data taking
- Neutral Particle Spectrometer and Compact Photon Source (Tanja)
- Magnet corrector coil sent to JLab - should go into TEDF in space in front of cleanroom
- Main coil should arrive end of July at JLab
- Two experiments (WACS and K_Long) being reviewed by PAC45
- Brief discussion of Compact Photon Source Concept and conceptual design status - more details in 6 July NPS meeting and special presentation at PAC45
- Gluex detectors (Greg)
- High magnetic field runs to test how magnetic field impacts sensors of the EIC DIRC detector planned for August at JLab
- Optimization of monitoring system for transporting Babar DIRC boxed to JLab ongoing - recent changes include cameras, mounting structures
- A 3-layer (3 refractive indices) EIC DIRC lens prototype has been ordered - test beam at CERN in August to test the optical performance
PBWO4 STUDIES AT CUA (Richard, Blessed, Casey)
- Discussion about prototype assembly
- mechanical assembly completed, calibration tests ongoing (vibration, rotation, weight sensors)
- thermal assembly ongoing
- Crystal characterization continuing
- light yield
- chemical analysis
- impact of crystal surface nonuniformity on resolution/detector performance
DETECTORS AT CERN (Rishabh)
- Discussion of an event that is tracked through the detectors including tracker, pixel
- track alignment in software ongoing - may be related to positioning of multiple detector layers
- Discussion about material budget analysis with CMS9 software
- red=phase 1 upgrade, blue=phase 0
DATA ANALYSIS AND SIMULATIONS (Andres, Salina)
- Discussion of CUA's kaon experiment in context of Hall C running over next year
- SHMS optics commissioning and further detector checkout
- Fall commissioning experiments may have data for initial calibration checks (need to check and negotiate kinematics)
- Discussion of software setup and aerogel detector data analysis from KPP run in preparation for fall running (Andres)
- Discussion of initial results from DVCS deadtime analysis (Salina)
NEXT MEETING: WED 5 JULY AT 10:00 AM (EDT)